OpenStudy (anonymous):

Several students decide to start a T-shirt company. After initial expenses of $280, they purchase each T-shirt wholesale for $3.99. They sell each T-shirt for $10.99. How many must they sell to break even? a. 10 shirts c. 30 shirts b. 20 shirts d. 40 shirts

OpenStudy (mathmale):

Mind if I start off by asking you to define "break even" in this context?

OpenStudy (anonymous):

like breaking the whole in half i guess i confused about it thats why im asking

OpenStudy (anonymous):

break even means when what you spend equals what you make Understand ?

OpenStudy (anonymous):

oh i get it now but can you still help me with the question?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I am not gonna take over mathmale's explaination

OpenStudy (mathmale):

Or, more precisely, one achieves "break even" when Revenue = Costs. We're talking about business principles here. What does "Revenue" mean from a business standpoint? Total costs?

OpenStudy (mathmale):

I do think it essential to know these terms before attempting to solve a "break even" problem.

OpenStudy (anonymous):

ok thank you

OpenStudy (mathmale):

Revenue: Amount of $ taken in by the shopkeeper when x units of product are sold at price p. Costs: Fixed costs plus variable costs = F + c*x, where F=fixed costs, c=unit cost, and x=number of units produced. Have you seen these definitions before?

OpenStudy (mathmale):

ggiirrll: What's your strategy for learning new vocabulary? We all need such strategies. Break even implies that Total Costs = Revenue. Total Costs = F + c*x = p*x Your job is to identify values of F, c and p and then solve this equation for x. Would you try that, please?